A man 5 feet tall casts a shadow of 5 feet on the ground. Find the angle of elevation of the sun ?

1
Expert's answer
2022-02-14T13:56:20-0500

"\alpha - \text{Angle of elevation}"

"h = 5ft \text{ male height}"

"h = 5ft \text{ shadow length}"

"\tan \alpha=\frac{h}{l}=\frac{5}{5}=1"

"\alpha =\arctan(1) = 45\degree"


"\text{Answer: }45\degree"